OpenStudy (anonymous):

find the equation of a line passing through (1,2) and (-2,5)

OpenStudy (anonymous):

solve the slope slope = (2-5) / (1 + 2) slope = -3/3 = -1 y = mx + b 2 = -1(1) + b 3 = b y = mx + b y = -x + 3 x + y = 3

OpenStudy (anonymous):

thank you